OpenStudy (anonymous):

Rewrite using a negative exponent. 5/x^2

OpenStudy (rock_mit182):

\[\frac{ 5 }{ x ^{2} }= 5*x ^{-2} \]

OpenStudy (anonymous):

what about 1/3^2

OpenStudy (rock_mit182):

you have to apply the same property: \[\frac{ 1 }{ a ^{n} }= 1*a ^{-n}\]

OpenStudy (rock_mit182):

@rabbithunter96 Easy isn't ?